OpenStudy (anonymous):

Which two square roots are used to estimate 5

OpenStudy (jhannybean):

estimate 5 or \(\sqrt{5}\)?

OpenStudy (jhannybean):

Because you can approx the square root values of prime numbers by comparing perfect squares, i.e 2^2 = 4 and 3^2 = 9